OpenStudy (anonymous):

@openstudier101x @MrNood Jenna has a bag that contains 7 red marbles and 25 blue marbles. She selects a marble at random, and then, without replacing the first one, selects another marble at random. What is the probability that Jenna selects a blue marble and then a red marble? Round your answer to the nearest percent. P(blue and red) ≈ %

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Usual without replacement stuff. As I've no recollection of you ever asking such a question, I'll give you one free one. 7 red 25 blue 32 total p(red on first draw) = 7/32 p(blue on first draw) = 25/32 The problem is about half done. Think about what happens after we removes one blue one.